Grade 1210 grade science

How much energy is given to each coulomb of charge passing through a 6 V battery?

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student


Amount of charge = 1 coulomb
Potential difference = 6V
Energy or work done = unknown
Work done or energy = V × Q[Q is the amount of charge flowing between two points at potential difference V = 6]
i.e., W = V × Q
W = 6 × 1
Work done or energy = 6 joules.

Thanks